13. A partile is in simple harmonic motion along the x axis. The amplitude of the motion is When it is at , its kinetic energy is K 5J and its potential energy (measured with 0 isU-3J. When it is at r -12 , the kinctic and potential energies are 0 at x CK-8J and U 0 K-0J and U-S sot Sure 14. The system of pulleys and a hanging weight shown to the right is in equilibrium due to the pulling force, P. If the weight, W, is 24N, the pulling force, P must be: a) 24NB) 12N C)N D) 4N E) 3N 15. To measure the mass of a planet with the same radius as Earth, an astronaut drops an object from rest (relative to the planct) from an altitude of one radius above the surface. When the object hits its speed is 4 times what it would be if the same experiment were carricd out for Earth. In units of Me (the mass of the Earth), the mass of the planct is B) 4 ME C) 8 ME p D) 16 ME E) 32 ME 3 16. A 5.0 m weightless strut, hinged to a wall, is uscd to support an 800-N block as shown. The horizontal and vertical components of the force of the hinge on the strut are: FH- 600 N, Fy 800 N Tf T-80 C)-Fn D) Fu 1200 N, Fy 800 N E) Fr 800 N, Fr-800 N 800 1 hi nge 17. The 600-N ball shown is suspended on a string AB and rests against the frictionless vertical wall. The string makes an angle of 30 with the wall. The linc AB extends through the center of the ball, and the contact point with the wall is at the same vertical height as the center of the ball. The magnitude of the tension in the string is A) 1200 N B) 690N C) 600N D) 520N (E) 300 N 600 N
Answer #1

13] Maximum U = total energy = 5 + 3 = 8 J

U = 0.5 k (x1)^2 =  0.5 k (0.5xm)^2 = 0.5 k xm^2 /4 = 8/4 = 2 J

K = 8 -2 = 6 J

option A is correct